What is the vertex of the quadratic function x^2+2x+2?
Answer by stanbon(75887)

vertex of the quadratic function x^2+2x+2
----------------------
Rewrite as y=x^2+2x+2
Complete the square as follows:
y-2+?=(x^2+2x+?)
y-2+1=(x^2+2x+1)
y-1=(x+1)^2
This follows the form y-k=(x-h)^2 where (h,k) is the vertex.
Your vertex is at h=-1, k=1
Cheers,
Stan H.